You are here: Foswiki>Commissioning Web>Logs>07Mar013_Log (08 Mar 2013, JuanGuerra)Edit Attach

UT 20130307 AO#2 SX

Observer:JCGuerra(LBTO),JChristou(LBTO)
Telescope Operator: David(LBTO)
SW Support: Cox,D (Tucson)
AO Support: JChristou(LBTO)
Telescope Support: Jeff U(LBTO)
Instrument Support: Tim S (LBTO)

Summary

RR installed in right side (RFBG). Check of the internal source with the CCD47. Test of the software patch to correct the new telescope rotator position (shift 180 degrees).

Detail

Log of the telescope and pupil re-rotator test:

7 March 2013

AGW2 Rotator Tests (R = Rotator angle)

19:08 check alignment and set-up R = 270 degrees
  1. :13 - Closed Loop with 10 modes dust seen on WFS camera due to beamsplitter - moving the q-stage moved too far and lost IRTC IRTC spot position @ 165,128 Appears to be a TT problem on 1-axis on BCU47.

19:44 Reboot the WFS 19:49 Came up in funny state - the BCU39 view was odd. Changed binning and it's now working okay. 19:53 Error in slopes computed Severely aberrated image Changing Z-stage Dust seen on WFS TT appears to be okay though Closed - optimized gain at 400 modes.

VERIFIED AO OPERATION CHANGING TO R=0 degrees

20:06 R = 0 degrees Pupil Re-rotator is still going Started to derotate again away from zero (The telescope rotator is fixed). Went to 360 and now going back down again. Alfio thinks it's a problem with the limit switch at 0 degrees - suggestion to go to -10 degrees

20:18 R=-10 degrees Pupil derotator goes to -5 degrees and stays. Closing Loop - 10 modes - Okay
  1. modes did not work? Open loop - recentering pupil
  2. modes again - good
  3. modes - good
  4. modes - good - some vignetting - okay
  5. modes - some shadow of dust is seen but good. (20:29) Taking data in different beam-splitter positions to verify if it's dust shadow.

20:34 R = -45 degrees Pupil derotator to 22.500 degrees Realigning the beam
  1. modes - good
  2. modes - good
  3. modes - good
  4. modes - good optimized gain - good - taking closed PS data

20:46 R = -80 degrees Pupil derotator to 40 degrees
  1. modes - good
  2. modes - good
  3. modes - good
  4. modes - good optimizing gain - good take closed loops data
20:52 IRTC Strehl measurement tests Focus IRTC problems getting it to focus because of aberrations

skipping the focussing stage here as we have to verify the patch works for the rotations

21:16 R = +60 degrees Pupil derotator to 330 degrees
  1. modes - good
  2. modes - good
  3. modes - good
  4. modes - some apparent rotation due to the match? open loop and close
  5. modes - good
  6. modes - good
  7. modes - good optimized gains - 400 modes drop gain of 0.1 in high-order to prevent oscillation.

21:28 R = +120 degrees Pupil derotates to 300 degrees trouble finding WFS illumination went back to ... R=80 degrees
  1. modes - good rotate to 120 - kept WFS loop okay
  2. modes - good
  3. modes - good
  4. modes - good
  5. modes - good optimize gains - good
  6. modes

21:46 R =+180 degrees Pupil derotates to 270 degrees
  1. modes - good
  2. modes - good
  3. modes - good
  4. modes - good optimized gain - good
  5. modes

21:54 R=+240 degrees Pupil derotates to 240 degrees
  1. modes - good
  2. modes - good
  3. modes - good
  4. modes - good optimized gain - good
  5. modes

22:00 R=+300 degrees Pupil derotates to 210 degrees
  1. modes - good
  2. modes - good
  3. modes - good
  4. modes - good optimized gain - good
  5. modes

22:07 R=+359 degrees Pupil derotates to 180.5 degrees
  1. modes - good
  2. modes - good
  3. modes - good
  4. modes - good optimized gain - good
  5. modes

22:15 R=+420 degrees Pupil derotates to 150 degrees
  1. modes - good
  2. modes - good
  3. modes - good
  4. modes - good optimized gain - good
  5. modes

22:30 Alfio rejoins as we are checking the o degree problem and the Pupil Derotator goes into a strange mode. Alfio finds a limit set in the software and we shall retry going to 0 degrees again.

Checked out the -10 degree and now it goes 365 degrees

-- JuanGuerra - 07 Mar 2013
Topic revision: r2 - 08 Mar 2013, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback